OpenStudy (anonymous):

Juan is 8 years older than his sister. In 3 years, he will be twice as old as she will be then. How old are they now?

OpenStudy (anonymous):

j = s + 8 j + 3 = 2(s+3) s+8+3 = 2s + 6 s+11 = 2s +6 s+11-6 = 2s s+5 = 2s 5 = 2s-s 5 = s s = 5 j = 5+8 5 and 13

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

Your Welcome!! :)